# Tunable non-Markovian dynamics in a collision model: an application to coherent transport

### v2

@@ IV.2 Coherent transport
Overall, we can conclude that non-Markovianity improves the coherent transport for strong system-environment couplings but hinders it if the coupling is weak. In the weak coupling regime, a Markovian environment enhances both the amount of coherence and the coherence time of the excitation. Note that these two mechanisms may not be mutually exclusive. For example, in the Fenna-Matthews-Olson photosynthetic complex [40], the electronic degrees of freedom of the bacteriochlorophyll molecules (the chain in our model) can couple both to the phonons of the molecular environment (non-Markovian) and to an external thermal bath (Markovian environment). Our results might capture both of these situations, suggesting that the coexistence of a weakly coupled thermal bath and a strongly coupled non-Markovian environment might determine an optimal condition for coherent excitonic transport.
V Conclusions
@@
In this work, we have proposed a collision model that allows direct control of the environmental degree of non-Markovianity. We have fixed the number and geometry of the reservoir qubits and induced information loss by applying depolarising channels on them. We have characterised the effect of the depolarising channel by measuring the degree of non-Markovianity of the model, showing how the intensity of the former directly relates to the latter.
@@
We have applied our model to study the quantum coherent excitonic transport on a chain of interacting qubits coupled to an environment. Our model permits the exploration of how different regimes of non-Markovianity and system-environment coupling affect coherent transport. Studying this aspect is important, for instance, in photosynthetic complexes, where the electronic degrees of freedom responsible for the coherent transport of an exciton can couple to both Markovian (e.g. solvent) and non-Markovian (phonons) environments. Our results show that a non-Markovian environment aids coherent transport only in the strong system-environment coupling regime, while a Markovian environment is surprisingly preferable in the weak coupling regime.
@@
This work provides some information-theoretic insights into how non-Markovianity affects coherent excitonic transport, avoiding the complications that arise from accurately modelling a real complex physical system. The system-independent characteristic of the model might make it suitable for applications to any quantum processes whose understanding involves the exchange of information between the system and the environment. The model could also be easily adapted to study more specific systems. For example, one could expand the dimensions of the chain and the environment, use more realistic Hamiltonians, or use two different environments, one Markovian and one non-Markovian. These generalisations might be relevant to inform the development of novel technological applications. The model could also be implemented on current quantum computers [41]. We leave these investigations for future work.
@@ Acknowledgments.
We thank Maria Violaris, Vlatko Vedral and Chiara Marletto for their sharp comments and fruitful discussions on this manuscript. S.R. thanks the Fondazione “Angelo Della Riccia”. G.D.P. thanks the Clarendon Fund and the Oxford-Thatcher Graduate Scholarship for supporting this research.

### v3

@@ III Application to Coherent transport
In conclusion, when the chain is the most efficient medium for the propagation of the excitation, i.e. $J_{res}/J_{chain}\ll 1$, and when $0.25\leq\Omega\leq 1$, we have seen that non-Markovianity improves the coherence of the transport for strong system-environment couplings but hinders it if the coupling is weak. Note that these two mechanisms may not be mutually exclusive. For example, in the Fenna-Matthews-Olson photosynthetic complex [40], the electronic degrees of freedom of the bacteriochlorophyll molecules (the chain in our model) can couple both to the phonons of the molecular environment (non-Markovian) and to an external thermal bath (Markovian environment). Our results might capture both of these situations, suggesting that the coexistence of a weakly coupled thermal bath and a strongly coupled non-Markovian environment might determine an optimal condition for coherent information transport.
(a) Population of the first system qubit $(\eta=0.4$)
(b) Population of the first system qubit $(\eta=1.2$)
(c) Population of the third system qubit $(\eta=0.4$)
(d) Population of the third system qubit $(\eta=1.2)$
Figure 7: Population elements $|\rho_{C}^{1,1}(k)|$ and $|\rho_{C}^{3,3}(k)|$ over multiple iterations of the protocol for different values of $\Omega$.
(a) $|\rho_{C}^{1,1}(k)|$-$|\rho_{C}^{1,1}(k)|_{NT}$ for $\eta=0.4$
(b) $|\rho_{C}^{1,1}(k)|$-$|\rho_{C}^{1,1}(k)|_{NT}$ for $\eta=1.2$
(c) $|\rho_{C}^{3,3}(k)|$-$|\rho_{C}^{3,3}(k)|_{NT}$ for $\eta=0.4$
(d) $|\rho_{C}^{3,3}(k)|$-$|\rho_{C}^{3,3}(k)|_{NT}$ for $\eta=1.2$
Figure 8: Difference between the populations $|\rho_{C}^{N,N}(k)|$ in the case $J_{chain}=10$, $J_{res}=1$ and the populations $|\rho_{C}^{N,N}(k)|_{NT}$ in the case $J_{chain}=J_{res}=0$ (no transport). The difference is plotted for the first and third qubits of the chain over multiple iterations of the protocol, for $\eta=0.4$ and $\eta=1.2$, and with different values of $\Omega$. The vertical axis has been truncated to improve the visibility of the graphs.
To complete the study, in Fig. 7 we plot $|\rho_{C}^{1,1}(k)|$ and $|\rho_{C}^{3,3}(k)|$ as a function of the number of iterations of the protocol $k$ for different values of $\Omega$, once again in the weak ($\eta=0.4$, Figs. 7a-7c) and strong ($\eta=1.2$, Figs. 7b-7d) coupling regimes previously identified, with $J_{chain}=10$ and $J_{res}=1$. We observe that the populations tend to $1/8$ unless $\Omega=0$, which corresponds to a fully unitary process and so does not have a steady state. The steady state for $\Omega\neq 0$ is the maximally mixed state $\mathds{I}^{\otimes 3}/8$ determined by the repeated application of the depolarising channel. In our model, the populations for $\Omega\neq 0$ are not indicative of the location of the excitation. Instead, taken together with the coherences, they show that the information initially located on the first qubit is gradually lost. This is not a limiting aspect of our model; it just reflects its focus on the information dynamics of the process. To get specific information about the location of the excitation, non-Markovianity should be implemented with a different model (see, for instance, [41]). Nevertheless, we expect that the excitation dynamics will also affect population changes during the initial iterations of the protocol.
To quantify this effect, in Fig. 8 we plot the differences between the populations $|\rho_{C}^{1,1}(k)|$ and $|\rho_{C}^{3,3}(k)|$ of Fig. 7 ($J_{chain}=10$, $J_{res}=1$), and the populations $|\rho_{C}^{1,1}(k)|_{NT}$ and $|\rho_{C}^{3,3}(k)|_{NT}$ in the case $J_{chain}=J_{res}=0$, i.e. when there is no transfer of information along the chain or the reservoir and so the change in populations is entirely due to information loss. We interpret these population differences as an indicator of the population change caused solely by excitation transport, excluding the effects of information loss. We immediately note that—apart from the case $\Omega=0$—the values of Fig 8 are much smaller than the corresponding values in Fig. 7, meaning that information loss is the dominant effect on population changes. Looking at Fig. 7 with this in mind, we observe that higher values of $\Omega$ correspond to a slower information loss in the weak coupling regime, signalled by a slower decay in the population of the first qubit (Fig. 7a) and a faster increase in the population of the third qubit (Fig. 7c). Conversely, we see that higher values of $\Omega$ determine a quicker information loss in the strong coupling regime (Figs. 7b-7d). This agrees with the results of Sec. II, Fig. 3a: for $0.25\leq\Omega\leq 1$, higher Markovianity hampers information loss in the weak coupling regime and accelerates it in the strong coupling regime.
Let us now focus on the effect of the information transport on the population changes as quantified by Fig. 8. In Figs. 8a-8b, negative values mean that the depopulation of the first qubit is quicker when there is transport on the chain and the reservoir ($J_{chain}\neq 0,J_{res}\neq 0$) compared to the case without transport ($J_{chain}=J_{res}=0$). Similarly, positive values in Figs. 8c-8d indicate that the transport contributes to the increase in the population. Overall, Fig. 8 points to an information transfer from the first to the third qubit before the information is lost in the long run (unless $\Omega=0$ for which there is no information loss). Moreover, we note that, for $0.25\leq\Omega\leq 1$, higher Markovianity generally enhances information transfer in the weak coupling regime and lowers it in the strong coupling regime. This is in line with our results on the coherence of the transport (Fig. 5), meaning that this effect holds for both the amount of information transferred and its coherence.
IV Conclusions
@@
In this work, we have proposed a collision model that allows direct control of the environmental degree of non-Markovianity. We have fixed the number and geometry of the reservoir qubits and induced information loss by applying depolarising channels on them. We have characterised the effect of the depolarising channel by measuring the degree of non-Markovianity of the model, showing how the intensity of the former directly relates to the latter. We have also studied how the information that the system loses is related to the degree of non-Markovianity and found a regime where, surprisingly, Markovianity reduces information loss.
@@
We have applied our model to study the quantum coherent information transport on a chain of interacting qubits coupled to an environment. Our model permits the exploration of how different regimes of non-Markovianity and system-environment coupling affect coherent transport. Studying this aspect is important, for instance, in photosynthetic complexes, where the electronic degrees of freedom responsible for the coherent transport of an exciton can couple to both Markovian (e.g. solvent) and non-Markovian (phonons) environments. Our results show that a non-Markovian environment aids the coherent transport only in the strong system-environment coupling regime. If the system is weakly coupled to the environment, we have observed a regime where Markovianity enhances the coherence of the transport, consistently with our results about information loss.
@@
This work provides some information-theoretic insights into how non-Markovianity affects coherent information transport, avoiding the complications that arise from accurately modelling a real complex physical system. The system-independent characteristic of the model might make it suitable for applications to any quantum processes whose understanding involves the exchange of information between the system and the environment. The model could also be easily adapted to study more specific systems. For example, one could expand the dimensions of the chain and the environment, use more realistic Hamiltonians, or use two different environments, one Markovian and one non-Markovian. These generalisations might be relevant to inform the development of novel technological applications. The model could also be implemented on current quantum computers [42]. We leave these investigations for future work.
@@ Acknowledgments.
We thank Maria Violaris, Vlatko Vedral, Chiara Marletto and an anonymous Referee for their sharp comments and fruitful discussions on this manuscript. S.R. thanks the Fondazione “Angelo Della Riccia”. G.D.P. thanks the Clarendon Fund and the Oxford-Thatcher Graduate Scholarship for supporting this research.